36x^2+25y^2+288x+200y+76=0 is the equation of a conic. Write the standard form of that conic.
Answer by mananth(16946)

36x^2+25y^2+288x+200y+76=0

complete the square for both x & y
36x^2 +288x + ------- +25y^2+200y+_______ = -76
36x^2 +288x + 576 +25y^2+200y+ 400 = -76 +576+400

36x^2 +288x + 576 +25y^2+200y+ 400 = 900
36(x+4)^2+25(y+4)^2=900
divide by 900
(x+4)^2/25 +(y+4)^2/36 =1